# Mitigating Label Noise using Prompt-Based Hyperbolic Meta-Learning in Open-Set Domain Generalization

### v1

@@ 5.3 Analysis of the Module Ablations
The ablation study of the individual components in our HyProMeta framework is presented in Table 13. This study is conducted on the PACS dataset across all label noise settings, with ResNet18 [24] used as the feature learning backbone. Two HyProMeta variants are considered in this analysis: w/o HYB-Meta and w/o NCA-Prompt. The w/o HYB-Meta variant incorporates only the new category-aware learnable prompt into the training pipeline, whereas the w/o NCA-Prompt variant utilizes only the hyperbolic prototype-based label noise-agnostic meta-learning.
@@
Overall, the results highlight the complementary roles of NCA-Prompt and HYB-Meta, where their integration in HyProMeta achieves superior performance. The NCA-Prompt contribute to the generalizability of embeddings, thereby augmenting the effectiveness of HYB-Meta in handling label noise in OSDG-NL task.
5.4 Analysis of t-SNE Visualizations
We provide t-SNE [62] visualization results on the PACS dataset, comparing our proposed approach with three baseline methods: BadLabel [75], MLDG [32], and MEDIC [66]. The visualizations of the learned embeddings for the test domain Photo are shown in Figure 3(a) to Figure 3(d), while those for the test domain Art Painting are provided in Figure 3(e) to Figure 3(h).
@@ 5.4 Analysis of t-SNE Visualizations
5.5 Analysis of the Confidence Score
@@
Our observations indicate that the confidence score distributions obtained by HyProMeta exhibit greater separation between seen and unseen categories. This separation highlights a key advantage of our proposed approach, as it enables the model to achieve a strong awareness of unseen categories and to assign them low confidence scores in the target domain. These findings further demonstrate the effectiveness of HyProMeta in handling the complexities of OSDG-NL task.
@@ 5.6 Ablation of the NCA-Prompt
We compare our proposed NCA-Prompt approach with several alternative variants: Asynchronous Opt., Adversarial Opt., and Fixed Crop.. These comparisons are presented in Table 14, with experiments conducted on the ResNet18 [24] backbone using the PACS dataset under a symmetric label noise ratio of $50\%$.
@@
(a) BadLabel [75]
(b) MLDG [56]
(c) MEDIC [66]
(d) HyProMeta (Ours)
(e) BadLabel [75]
(f) MLDG [56]
(g) MEDIC [66]
(h) HyProMeta (Ours)
Figure 4: Confidence score visualization of learned representations on PACS using ResNet18 [24] under symmetric label noise with ratio $50\%$ and asymmetric label noise with ratio $50\%$.
@@ 5.7 Ablation of HYB-Meta
The ablation study of the HYB-Meta approach is presented in Table 15. This study is conducted using ResNet18 [24] as the backbone on the PACS dataset under a label noise ratio of $50\%$. Three variants are examined in this analysis, specifically w/o label corr., w/o cross domain, and w/ euc. prototype.
w/o label corr. refers to the removal of the hyperbolic prototype-based label correction mechanism in HYB-Meta. Without this correction, our approach demonstrates a $2.53\%$ improvement in OSCR, highlighting the critical role of label correction in the meta-learning process for the OSDG-NL task. w/o cross domain indicates that both the meta-train and meta-test partitions are derived from the same source domain, rather than incorporating cross-domain data. Under this setting, HYB-Meta achieves a $4.98\%$ OSCR improvement, emphasizing the benefits of cross-domain data partitioning in enhancing model performance for the OSDG-NL task. w/ euc. prototype involves replacing the categorical prototypes computed in hyperbolic space with those computed in Euclidean space. This substitution results in a $3.16\%$ OSCR performance drop, demonstrating the effectiveness of hyperbolic categorical prototypes in leveraging the unique properties of hyperbolic space for OSDG-NL.
@@ 5.8 Ablation Regarding the Step Length of the Prototype Calculation
Table 16: Ablation study of $N_{epoch}$ on PACS with ResNet18 [24] backbone, where Photo is selected as the test domain.
@@ 6 Conclusion
Future works: In the future we will explore the OSDG-NL task in other down-stream tasks, e.g., video classification, semantic segmentation, and multi-modal learning.
@@ Acknowledgment
The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation) with a start date of April 1, 2024. This work was also partially supported by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. The authors also acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. This project was also partially supported by the National Natural Science Foundation of China under Grant No. 62473139.

### v2

@@ 5.7 Analysis on the Module Ablations
The ablation study of the individual components in our HyProMeta framework is presented in Table 13. This study is conducted on the PACS dataset across all label noise settings, with ResNet18 [he2016deep] used as the feature learning backbone. Two HyProMeta variants are considered in this analysis: w/o HYB-Meta and w/o NCA-Prompt. The w/o HYB-Meta variant incorporates only the new category-aware learnable prompt into the training pipeline, whereas the w/o NCA-Prompt variant utilizes only the hyperbolic prototype-based label noise-agnostic meta-learning.
@@
Overall, the results highlight the complementary roles of NCA-Prompt and HYB-Meta, where their integration in HyProMeta achieves superior performance. The NCA-Prompt contributes to the generalizability of embeddings, thereby augmenting the effectiveness of HYB-Meta in handling label noise in the OSDG-NL task.
5.8 Analysis on t-SNE Visualizations
We provide t-SNE [van2008visualizing] visualization results on the PACS dataset, comparing our proposed approach with three baseline methods: BadLabel [zhang2024badlabel], MLDG [li2018learning], and MEDIC [wang2023generalizable]. The visualizations of the learned embeddings for the test domain Photo are shown in Figure 3(a) to Figure 3(d), while those for the test domain Art Painting are provided in Figure 3(e) to Figure 3(h).
@@ 5.8 Analysis on t-SNE Visualizations
5.9 Analysis on the Confidence Score
@@
Our observations indicate that the confidence score distributions obtained by HyProMeta exhibit greater separation between seen and unseen categories. This separation highlights a key advantage of our proposed approach, as it enables the model to achieve a strong awareness of unseen categories and to assign them low confidence scores in the target domain. These findings further demonstrate the effectiveness of HyProMeta in handling the complexities of the OSDG-NL task.
@@ 5.10 Ablation of the NCA-Prompt
We compare our proposed NCA-Prompt approach with several alternative variants: Asynchronous Opt., Adversarial Opt., and Fixed Crop.. These comparisons are presented in Table 14, with experiments conducted on the ResNet18 [he2016deep] backbone using the PACS dataset under a symmetric label noise ratio of $50\%$.
@@
(a) BadLabel [zhang2024badlabel], D:Photo, LNSR:50% sym
(b) MLDG [shu2019meta], D:Photo, LNSR:50% sym
(c) MEDIC [wang2023generalizable], D:Photo, LNSR:50% sym
(d) HyProMeta (Ours), D:Photo, LNSR:50% sym
(e) BadLabel [zhang2024badlabel], D:Photo, LNSR:50% asym
(f) MLDG [shu2019meta], D:Photo, LNSR:50% asym
(g) MEDIC [wang2023generalizable], D:Photo, LNSR:50% asym
(h) HyProMeta (Ours), D:Photo, LNSR:50% asym
Figure 4: Confidence score visualization of learned representations on PACS using ResNet18 [he2016deep] under symmetric label noise with a ratio of $50\%$ and asymmetric label noise with a ratio of $50\%$. D indicates the target domain and LNSR indicates the label noise setting.
@@ 5.11 Ablation of HYB-Meta
The ablation study of the HYB-Meta approach is presented in Table 15. This study is conducted using ResNet18 [he2016deep] as the backbone on the PACS dataset under a label noise ratio of $50\%$. Three variants are examined in this analysis, specifically w/o label corr., w/o cross domain, and w/ euc. prototype.
w/o label corr. refers to the removal of the hyperbolic prototype-based label correction mechanism in HYB-Meta. Without this correction, our approach demonstrates a $2.53\%$ improvement in OSCR, highlighting the critical role of label correction in the meta-learning process for the OSDG-NL task. w/o cross domain indicates that both the meta-train and meta-test partitions are derived from the same source domain, rather than incorporating cross-domain data. Under this setting, HYB-Meta achieves a $4.98\%$ OSCR improvement, emphasizing the benefits of cross-domain data partitioning in enhancing model performance for the OSDG-NL task. w/ euc. prototype involves replacing the category prototypes computed in hyperbolic space with those computed in Euclidean space. This substitution results in a $3.16\%$ OSCR performance drop, demonstrating the effectiveness of hyperbolic category prototypes in leveraging the unique properties of hyperbolic space for OSDG-NL.
@@ 5.12 Ablation Regarding the Step Length of the Prototype Calculation
Table 26: Ablation study of $N_{epoch}$ on PACS with ResNet18 [he2016deep] backbone, where Photo is selected as the test domain.
@@ 6 Conclusion
Future work: In the future, we will explore the OSDG-NL task in other downstream tasks, e.g., video classification, semantic segmentation, and multi-modal learning.
@@ Acknowledgment
The project served to prepare the SFB 1574 Circular Factory for the Perpetual Product (project ID: 471687386), approved by the German Research Foundation (DFG, German Research Foundation) with a start date of April 1, 2024. This work was also partially supported by the SmartAge project sponsored by the Carl Zeiss Stiftung (P2019-01-003; 2021-2026). This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. The authors also acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. This project was also supported partially by the National Natural Science Foundation of China under Grant No. 62473139, partially by the Hunan Provincial Research and Development Project under Grant No. 2025QK3019, and partially by the State Key Laboratory of Autonomous Intelligent Unmanned Systems (the opening project number ZZKF2025-2-10).